solve 2x+52x+3=x+3/x-1
Answer by kirti251(6)

2x+52x+3=x+3/x-1
(x-1)(2x+52+3)=x+3
x(2x+52+3)-1(2x+52+3)=x+3
2x+52x+3x-2x-52-3=x+3
55x-49=x+3
55x-x=49+3
54x=52
x=52/54
x=26/27